### v2

# Leveraging the true depth of LLMs

Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities at the cost of high compute requirements. Recent studies have demonstrated that intermediate layers in LLMs can be removed or reordered without substantial accuracy loss; however, this insight has not yet been exploited to improve inference efficiency. Leveraging observed layer independence, we propose a novel method that groups consecutive layers into pairs evaluated in parallel, effectively restructuring the computational graph to enhance parallelism. Without requiring retraining or fine-tuning, this approach achieves an inference throughput improvement of 1.05x–1.20x on standard benchmarks, retaining 95%–99% of the original model accuracy. Empirical results demonstrate the practicality of this method in significantly reducing inference cost for large-scale LLM deployment. Additionally, we demonstrate that modest performance degradation can be substantially mitigated through lightweight fine-tuning, further enhancing the method’s applicability.
1 Introduction
Figure 1: The effect of LP on execution time (4K tokens) and perplexity (measured against RedPajama[1]). The bigger Llama model achieves both better perplexity and faster execution speeds than the smaller Llama model.
The rapid advancement of LLMs has revolutionized Artificial Intelligence applications across industries. However, the ever-increasing computational demands of these models, with parameters often numbering hundreds of billions, present significant commercial challenges. Efficient inference is crucial for organizations that deploy these models at scale, as it directly impacts operational costs, user experience, and environmental sustainability [2, 3, 4]. Monthly cloud computing expenses for LLM inference can reach millions of dollars for high-traffic applications, making optimization techniques essential. In addition, reducing inference latency is critical for real-time applications and for deploying models on devices with more limited compute. Thus, the development and implementation of efficient inference methods has become a key differentiator in the competitive AI market, driving both innovation and profitability.
+FFNLNMHALN++FFN1LNFFN2LNMHA1LNMHA2LN+(a)(b)
Figure 2: Comparing a normal transformer block with our layer parallel implementation. In (a) we show a normal Transformer Layer. In (b) we illustrate our layer parallelization, where each column runs separately on one or multiple GPUs. The residuals are gathered and synchronized through a reduce operation across all GPUs. The LayerNorms’ weights are copied from the original blocks, and each divergent path uses the normalization from its original layer.
LLMs have evolved to incorporate architectures with hundreds of layers [5, 6]. These models are constructed from stacked transformer blocks, each comprising attention and feedforward subblocks, with a residual stream traversing the entire architecture to facilitate efficient gradient backpropagation during training. This architectural choice parallels the design principles of ResNets [7], where research has shown that network depth may be partially redundant, allowing layer reordering without significant performance degradation [8]. Recent investigations have revealed similar flexibility in transformer architectures [9], where interventions such as layer removal and swapping are applied without large performance degradations. Although these findings challenge our understanding of LLMs’ true effective depth, their potential for optimizing inference efficiency remains unexplored.
Stemming from this line of work, we set to find approaches that reduce the depth of LLMs. We do so by exploring several interventions to the computational graph of pre-trained LLMs such as shuffling, pruning or merging layers. As a result of this search, we find that multiple consecutive block pairs can be processed in parallel while maintaining acceptable performance across perplexity and In-Context Learning (ICL) benchmarks. We introduce a novel parallelism approach, Layer Parallelism (LP), that enhances inference speed when making use of tensor parallelism with a minimal drop in performance. Furthermore, we show that the performance degradation observed across various benchmarks can be partially mitigated through targeted fine-tuning procedures.
Contributions. Our contributions can be summarized as follows:
  • We explore the space of interventions on a pre-trained LLM layers, and find that some transformations, such as contiguous parallelization, preserve model performance
  • We find that we can define a parallelization transform on the computational graph of two sequential Transformer layers, and stack this parallelization operation to several sequential pairs of layers without loosing significant ICL performance. Our approach, which we call Layer Parallelism (LP), can be applied to existing Transformer models and does not require retraining.
  • We show that by fine-tuning the LP blocks we can recover some of the lost performance, while retaining the previously obtained speed-up.
As a result, the rearrangement of the computational graph allows for faster inference speeds when serving the LLM using tensor parallelism. The speed-up mainly arises from the decrease in the number of all-reduce operations between GPUs. This is because LP uses two all-reduce operations per pair of transformer decoder layers, while normal layers would require four all-reduce operations.
2 Related work
The effective depth of Deep Networks. Theoretically, given enough width, any feed-forward network with at least one hidden layer can model any function [10]. In practice, it is easier to achieve high expressivity by increasing the model’s depth. However, naively increasing network depth can complicate optimization, since the gradients now have to flow through many layers. To alleviate this problem, ResNets [7] introduced skip connections at regular intervals to allow an easy flow of the gradient to the first layers. Alternatively, Inception [11] explored approaches to boost computational power by adding additional processing units along different parallel pathways in the computational network, rather than just along a single sequential path. A unification of both methods can be found in the Highway Networks [12], where the skip connection of the residual blocks consists of another block of compute. Nowadays, residual connections are ubiquitous in large models.
(a) Shuffling.
(b) Pruning.
(c) Merging.
(d) Parallel.
(e) 2-Parallel.
Figure 3: Changes in perplexity when applying transformations on contiguous stretches of layers. Each one of the five heatmaps above correspond to a transformation of a group of consecutive layer, where the row index $s$ corresponds to the first layer of the group, and the column index $e$ to the last. The color coding indicates how the perplexity—estimated on a subset of RedPajama [1]—is impacted by the corresponding modification of the model. The perplexity for the base Llama 2 7B model is $6.2$. In (a), we shuffle—for each forward—the layers from $s$ to $e$. We can see that many consecutive layers can be shuffled with little impact on the overall perplexity. For instance, shuffling layers $15$ to 2525252510101010 layers in total—raises the perplexity only to $9.1$. In (b), we prune contiguous stretches of layers. We can see that not many blocks can be removed without starting to significantly degrade the perplexity. In (c) we merge contiguous layers. The results with merging are near identical to those for pruning. This reveals there is no advantage in merging layers, most likely a results of averaging matrices not originating from the same initial values. In (d) we run contiguous blocks in parallel. Given the success of shuffling, it makes sense that this approach works well. Running blocks $17$ to $27$ raises the perplexity to $9.3$. Finally, in (e) we run pairs of consecutive layers in parallel. As a result, we can parallelize much longer stretches of layers. For instance, we can apply this transformation from layer $4$ to $29$ and only increase the perplexity to $9.1$. This reduces the depth of the model from $32$ to $19$. This result makes it possible for us to leverage this parallelism for faster inference as we discuss in § 4.
Efficient inference of LLMs. Several complementary approaches exist for enhancing the computational efficiency of large-scale models, primarily through pruning/sparsity, quantization, and parallelism. Pruning[13, 14, 15, 16] constitutes a dimensional reduction methodology that systematically eliminates redundant parameters while preserving model performance, thereby introducing architectural sparsity. This methodology is founded on empirical evidence demonstrating that neural networks frequently exhibit overparameterization, containing numerous weights with negligible contributions to the output. Through sophisticated pruning strategies, the inherent sparsity support in contemporary accelerators can be leveraged to enhance both memory utilization and computational efficiency [17, 18]. Early-exit[19, 20] can be seen as a way of runtime layer-wise pruning, which halts the LLM forward pass when the next token certainty is high in the intermediate layers. This approach can be also thought as a way of reducing the effective depth of the model at test time. In contrast, quantization encompasses the transformation of floating-point numerical representations (predominantly FP32) into reduced-precision integer formats, such as INT8 or INT4 [21, 15, 22]. When implemented on hardware accelerators, these lower-precision representations facilitate superior memory bandwidth utilization, addressing a primary bottleneck in modern large-scale models [23]; moreover, integer-based computations yield enhanced processing speed and substantially improved energy efficiency [24]. Finally, parallelization techniques during inference, such as tensor and pipeline parallelism, enable the distribution of computational workload across multiple accelerators, thereby reducing latency and increasing throughput, although this often requires careful consideration of communication overhead and load balancing [25, 26].
Parallel attention-feedforward fusion. GPT-J[27] introduced a parallel formulation of the transformer decoder layer, executing attention and feedforward sub-blocks concurrently:
$y=x+\text{MHA}(\text{LayerNorm}(x))+\text{MLP}(\text{LayerNorm}(x))$
For models trained with tensor parallelism, this modification halves the number of required all-reduce operations. Additionally, it reduces memory bandwidth usage by eliminating one read and write operation of hidden states from HBM. The input projections for the attention and MLP sub-blocks can also be fused into a single kernel, further increasing arithmetic density. Consequently, training time is reduced by approximately 15% without observable performance degradation. PaLM[28] also adopted this formulation, noting that negative effects from deviating from the standard transformer self-attention diminish with increasing model size. This parallel formulation continues to be employed in more recent LLMs, including Gemini 1.5 Flash[29].
We differ from these approaches by focusing specifically on already-trained models with the standard transformer formulation[6, 30, 31], rather than training a model from scratch. By rearranging the computational graph, we merge two consecutive transformer-decoder layers by first executing both attention subblocks and both feedforward subblocks sequentially (Figure 2). Similar to GPT-J and PaLM, the primary source of acceleration in our method arises from halving the number of all-reduce operations across the layers where we apply LP.
Parallelism via Computational Graph Optimization. Recent research has investigated architectural layer-level optimization strategies to enhance transformer model inference efficiency. The Staircase Transformer [32] implements parallel layer execution with dynamic recurrent computation based on model requirements. Similarly, the Staggering Transformer [33] achieves layer parallelization by connecting layer $l_{k}$ at time step $t$ to both the $(t-1)$ output of layer $l_{k-1}$ and the $t$ output of layer $l_{k-2}$. To the best of our knowledge, no research has addressed the fusion of consecutive layers through tensor parallelism.
3 Effective Depth
[…]
Figure 4: Diagram of transformations applied in § 3. Diagrams (a,b,c,d) represent respectively shuffling, merging, pruning and parallel.
In this section we investigate the effective depth of pretrained LLMs by applying several transformations and measuring the resulting perplexity degradation. We reveal loose dependencies between intermediary layers. The transformations consist of shuffling, merging, and pruning transformer layers. To avoid the combinatorial explosion resulting from considering all the possible subsets of transformer layers, we instead apply our transformations to all the contiguous stretch of layers. If $L=\{\ell_{1},\ldots,\ell_{N}\}$ are the ordered layers, then we apply our transformations to all the sub-list $\{\ell_{i}\}_{i=s}^{e}$ with $1\leq s\leq e\leq N$. Previous works have shown that—at least when considering pruning—the importance of layers is well behaved, with low importance layers close to one another [34], which justifies considering contiguous stretch of layers only.
Shuffling blocks. We start by shuffling contiguous stretches of layers, re-ordering the layers according to random permutations. Results are shown in Fig. 3.(a). While shuffling the early and two last layers significantly raises the perplexity, there are large stretches of blocks which can be shuffled with surprisingly low impact on the perplexity. For instance, one can shuffle the layers $\{\ell_{i}\}_{15\leq i<25}$ of Llama 2 7B and only have a increase in perplexity of $2.9$. This goes against the classical belief that models build deeply hierarchical representations, where features in previous layers are leveraged to build more complex features in later layers. We interpret this as multiple layers working at the same level of abstraction. Using these insights, we define the effective depth of an LLM as the shortest depth required to efficiently leverage existing latent representations without a significant loss of performance. This reveals an important level of layer decoupling within the model.
Running blocks in parallel. The observed layer decoupling suggests that specific transformer operations may be executed independently, providing an opportunity for parallel computation. More precisely, let’s consider two sequential transformer layers $\ell_{k}$ and $\ell_{k+1}$, each comprising attention and Feed-Forward Network (FFN) sub-blocks ($A_{k}(\cdot)$ and $F_{k}(\cdot)$, respectively). The standard sequential output ${\bm{y}}$ for these layers, given an input ${\bm{x}}$, is given by:
${\bm{y}}={\bm{x}}$ $+\text{A}_{k}({\bm{x}})$
$+\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$
$+\text{A}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}})))$
$+\text{F}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$ (1)
$+\text{A}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))))$ (SEQ)
The highlighted terms represent the first block’s contribution to the second block’s processing. Given the observed layer independence, we can hypothesize that these terms have minimal impact, leading to the following approximation:
$\hat{{\bm{y}}}={\bm{x}}$ $+\text{A}_{k}({\bm{x}})+\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$ (2)
$+\text{A}_{k+1}({\bm{x}})+\text{F}_{k+1}({\bm{x}}+\text{A}_{k+1}({\bm{x}}))$ (PAR)
This approximation enables parallel execution of blocks $\ell_{k}$ and $\ell_{k+1}$ through divergent computational paths. We experiment with running contiguous stretches of layers in parallel and show our results in Fig. 3(d). We observe results similar to shuffling. Unlike shuffling, this approach allows us to potentially improve the runtime through enhanced parallelism. We show how we can, for instance, run layers $17$ to $27$ in parallel, only losing $3.1$ perplexity points, while reducing the depth of the model from $32$ to $23$.
Contiguous 2-parallel. Instead of parallelizing long stretches of layers, we experiment with running pairs of consecutive layers in parallel. This springs from the assumption that local ordering matters less than global ordering, i.e. shuffling consecutive layers introduces fewer potential issues than shuffling layers separated by larger distances. As an example, if we apply the proposed transformation to layers $\{\ell_{15},\ell_{16},\ell_{17},\ell_{18},\ell_{19}\}$, it would result in the following process: (1) the two layers $\{\ell_{15},\ell_{16}\}$ process the input in parallel (according to equation (PAR)), (2) the output is forwarded to layers $\{\ell_{17},\ell_{18}\}$ which process it in parallel; finally, in (3) their joint output is fed to layer $\ell_{19}$ which processes it on its own as any normal layer. The effect of such transformation of the compute graph on the perplexity can be seen in Fig. 3(e). Remarkably, it is possible to run wide stretches of consecutive pairs of blocks in parallel with only a minor degradation of perplexity. For instance, one can apply this transformation from layer $4$ to layer $29$ with only a degradation of perplexity of $2.9$, while reducing the model depth from $32$ to $19$. The success of this approach led us to also try running triplets of consecutive layers in parallel, but we found it to perform worse.
Exploring other transformations. We also experiment with pruning (Fig. 3(b)) and merging (Fig. 3(c)). Pruning has already been studied in several prior works [35, 36]. While pruning a layer leads to a larger perplexity increase compared to parallelizing two blocks, it offers the advantage of reducing the model’s parameter count, which results in improved throughput and more efficient memory usage.
4 Efficient Parallelization of Blocks
[…]
Figure 5: LP attention implementation. This diagram shows the implementation of the LP attention from Fig. 2b. The stacked layers represent different GPUs, the colors indicate different layers and the arrows express linear projections. In this case, the number of GPUs and the number of parallelized layers coincides and is two, which is the set-up that we use for all our experiments in this work.
If we naively tried to run two attention or MLP blocks by fusing their execution on a single GPU, we would not see any noticeable improvement on the inference speed. This is due to the intensive compute and memory bandwidth requirements of LLMs. For this reason we focus our attention on the tensor parallel setting, where each module’s weights are split over two or more GPUs. Rearranging the computational graph of two contiguous layers like in Fig. 2b effectively reduces the number of inter-GPU synchronizations by half. Now, each divergent path is computed in parallel over multiple accelerators, and only the single intermediate and final results need to be synchronized. While this approach is not numerically equivalent to (PAR), we nonetheless—and quite surprisingly— show that it works well on already trained models, circumventing the need to train from scratch.
LP Multi-Head Attention. Traditional tensor parallelism in MHA distributes attention heads evenly across GPUs[37], performing self-attention and output projection locally before gathering results through an all-reduce summation. Each GPU processes tensors of dimensions $Q,K,V,att\in\mathbb{R}^{T\times\frac{D}{g}}$, where $T$ is sequence length, $D$ is feature dimension, and $g$ is the number of parallel workers. The local output projection produces $o_{i}\in\mathbb{R}^{T\times D}$ for each worker $i$, with rank $\frac{D}{g}$, and then the all-reduce operation restores full rank. To implement LP, we increase the depth of the query, key, and value weight matrices ($W_{Q},W_{K},W_{V}\in\mathbb{R}^{(g_{n}\cdot h_{d})\times D}$) and widen the output projection ($W_{O}\in\mathbb{R}^{D\times(n_{h}\cdot h_{d})}$), where $n_{h}$ represents heads per GPU and $h_{d}$ is head dimensionality. The reduction operation now will simultaneously compute the full-rank output projections and the sum of all parallel layers (Fig. 5).
LP FFN. Standard tensor parallelism for single-hidden-layer FFNs splits the first layer’s output across devices, generates low-rank outputs from the second layer, and sums them through reduction. To parallelize two FFN layers, we double the first layer’s output dimensionality and perform separate output (low-rank) projections for each layer. A single reduction operation then serves the dual purpose of computing full outputs for each layer and combining their results, as shown in Fig.  2(b). In summary, LP for FFN just concatenates the up-projection weights and continues as normal TP, allowing for multiple GPUs to be allocated per parallelized layer.
Handling LN. Since we asume at least one GPU per parallelized layer, we can assign each original LN to the divergent path that contains the attention and FFN blocks from its original layer.
5 Experiments & Results
(a) Llama2 7B.
(b) Llama3.2 3B.
Figure 6: Perplexity when running pairs of consecutive layers in parallel. Perplexity of Llama2 7B and Llama3.2 3B models on the test set of RedPajama[1] when applying Layer Parallelism to $\Delta$ consecutive layers. The parallelized interval for each data point is [end indexΔ,end $[\text{end index}-\Delta,\text{end index}]$.
Table 1: 5-shot In-Context Learning accuracies across standard benchmarks. Effective Depth shows the minimum number of sequential operations from input to output. Par. Layers indicates the range of consecutive layers where pairs were processed in with Layer Parallelism. *ifeval was evaluated on 0-shot performance.
Eff. Depth MMLU PiQA Arc E. Arc C. WinoG OBQA hswag GSM8K ifeval* Avg.
Llama 2 7B
32 (Base) 0.4583 0.8009 0.8106 0.5196 0.7411 0.4520 0.7821 0.6521
27 (Ours) 0.4625 0.7933 0.8005 0.5094 0.7348 0.4600 0.7782 0.6484
26 (Ours) 0.4588 0.7927 0.7976 0.4983 0.7340 0.4460 0.7745 0.6431
25 (Ours) 0.4532 0.7851 0.7917 0.4949 0.7340 0.4440 0.7673 0.6386
24 (Ours) 0.4083 0.7845 0.7841 0.4839 0.7190 0.4360 0.7578 0.6248
23 (Ours) 0.3519 0.7829 0.7677 0.4488 0.6922 0.4240 0.7368 0.6006
Llama 3.2 3B
28 (Base) 0.5610 0.7992 0.7807 0.4872 0.7214 0.4520 0.7557 0.6510
24 (Ours) 0.5508 0.7856 0.7521 0.4753 0.7167 0.4420 0.7384 0.6373
23 (Ours) 0.5481 0.7748 0.7399 0.4735 0.7119 0.4200 0.7303 0.6284
22 (Ours) 0.4693 0.7666 0.7264 0.4497 0.6914 0.4180 0.7193 0.6058
21 (Ours) 0.3890 0.7519 0.6839 0.4061 0.6638 0.4020 0.6847 0.5690
20 (Ours) 0.3107 0.7416 0.6481 0.3652 0.6227 0.3620 0.6407 0.5273
Llama 2 7B (Chat)
32 (Base) 0.4727 0.7769 0.7963 0.4906 0.7206 0.3300 0.5887 0.2297 0.4365 0.5380
27 (Ours) 0.4746 0.7709 0.7803 0.4881 0.7182 0.3380 0.5839 0.2221 0.4736 0.5389
26 (Ours) 0.4724 0.7666 0.7748 0.4701 0.7151 0.3400 0.5779 0.1903 0.4856 0.5325
25 (Ours) 0.4767 0.7633 0.7719 0.4650 0.7009 0.3320 0.5724 0.1478 0.4796 0.5233
24 (Ours) 0.4547 0.7655 0.7567 0.4369 0.6788 0.3040 0.5582 0.0963 0.4149 0.4962
23 (Ours) 0.4271 0.7524 0.7428 0.4113 0.6654 0.3080 0.5406 0.0697 0.3765 0.4771
Llama 3.2 3B (Instruct)
28 (Base) 0.5956 0.7709 0.7929 0.4650 0.7024 0.3080 0.5253 0.6475 0.6715 0.6088
24 (Ours) 0.5914 0.7622 0.7660 0.4488 0.7009 0.3000 0.5131 0.4587 0.6031 0.5716
23 (Ours) 0.5905 0.7546 0.7618 0.4471 0.6898 0.2780 0.5116 0.3571 0.5995 0.5544
22 (Ours) 0.5612 0.7530 0.7504 0.4573 0.6677 0.2880 0.5088 0.1001 0.5564 0.5159
21 (Ours) 0.5072 0.7470 0.7256 0.4019 0.6409 0.2840 0.4947 0.0311 0.4904 0.4803
20 (Ours) 0.4070 0.7410 0.7075 0.3942 0.6275 0.2900 0.4755 0.0311 0.4448 0.4576
Qwen3 4B (Instruct)
36 (Base) 0.7016 0.7644 0.8476 0.5879 0.6630 0.3720 0.5277 0.8499 0.4209 0.6372
31 (Ours) 0.6887 0.7459 0.8182 0.5367 0.6598 0.3520 0.5010 0.5375 0.3549 0.5772
30 (Ours) 0.6749 0.7497 0.8184 0.5239 0.6511 0.3340 0.4895 0.3677 0.3813 0.5545
29 (Ours) 0.6356 0.7470 0.7988 0.5043 0.6322 0.3300 0.4725 0.1797 0.3657 0.5184
28 (Ours) 0.5395 0.7432 0.7929 0.4898 0.6275 0.3000 0.4524 0.1266 0.3465 0.4909
27 (Ours) 0.4409 0.7296 0.7668 0.4454 0.5967 0.2960 0.4322 0.0356 0.2782 0.4468
26 (Ours) 0.3795 0.7062 0.7231 0.4002 0.5675 0.2840 0.4064 0.0318 0.3165 0.4239
Qwen3 14B (Instruct)
40 (Base ) 0.7883 0.8090 0.8775 0.6621 0.7451 0.4040 0.6133 0.8226 0.4652 0.6875
35 (Ours) 0.7792 0.7982 0.8695 0.6468 0.7356 0.3880 0.5825 0.6164 0.4376 0.6504
34 (Ours) 0.7742 0.7889 0.8561 0.6271 0.7388 0.3780 0.5792 0.7536 0.4700 0.6629
33 (Ours) 0.7588 0.7905 0.8561 0.6058 0.7269 0.3740 0.5693 0.6983 0.4376 0.6464
32 (Ours) 0.7370 0.7889 0.8481 0.5973 0.7103 0.3680 0.5575 0.6467 0.4017 0.6284
31 (Ours) 0.7156 0.7856 0.8552 0.5998 0.6859 0.3640 0.5412 0.4822 0.3801 0.6011
30 (Ours) 0.6543 0.7758 0.8384 0.5589 0.6732 0.3460 0.5197 0.2343 0.3657 0.5518
29 (Ours) 0.5954 0.7688 0.8215 0.5256 0.6575 0.3260 0.4945 0.1585 0.3381 0.5207
Table 2: Benchmark accuracy restoration through finetuning on Qwen3 4B with an effective depth of 27.
Finetuning Steps MMLU Arc C. GSM-8K (%)
0 (Ours) 0.4409 0.4454 0.0356
1024 (Ours) 0.6063 0.4932 0.4011
4096 (Ours) 0.6196 0.5102 0.4829
8192 (Ours) 0.6134 0.5162 0.4943
Qwen3 4B (Base) 0.7014 0.5887 0.8514
Figure 7: Wall clock time to complete the following inference tasks: KV Cache pre-filling, autoregressive generation, and single token generation with a pre-filled KV Cache. $\Delta$ indicates how many layers have been merged using LP (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The gains in inference speed are directly proportional to the grade of LP. The 1-token generation task for Llama 3.2 3B does not saturate the GPU compute until a sequence length of 2048. Even in this regime, LP benefits from considerable speed-ups.
In this section, we evaluate Layer Parallelism across three dimensions: inference speed improvements, impact on In-Context Learning performance, and the potential to recover model accuracy through targeted fine-tuning of parallelized layers.
Experimental protocol. For all our experiments, we use a node with two A100 SXM4 80Gb GPUs, four AMD EPYC 7742 CPUs, and 512Gb of RAM. We test for varying sequence lengths, up to $4096$ (Llama’s context window), with a batch size of $1$ unless indicated otherwise. We consider two models of the Llama family: Llama2 7B, and Llama3.2 3B, as well as two sizes from Qwen3: 4B and 14B. Given a desired effective depth, we replace the required number of normal layers with LP layers. For Llama 2 7B and 3.2 3B, the LP layers are selected based on the configuration that minimized the PPL for a given amount of LP (Fig. 6). For Qwen3, LP is applied until the 4th to last decoder layer. The rest of the layers implement the tensor parallel approach as described in [37]. For evaluation, we measure the ICL 5-shot accuracies using the lm-eval package [38]. We test the ICL accuracy of the models on several tasks: MMLU [39], PiQA [40], ARC Easy, ARC Challenge, Winogrande [41], OpenBookQA [42], Hellaswag [43], GSM-8K [44] and ifeval [45]. The perplexity of the models is always evaluated against a subset of the test set of RedPajama [1].
Impact of LP on PPL and ICL accuracies. We first examine how perplexity evolves when applying LP across layer sequences of different lengths and depths. Fig.6 reveals a common optimal sequence end-index minimizing perplexity, found at layers 28 for Llama2 7B and 25 for Llama3.2 3B. Table1 compares the In-Context Learning performance across models with varying effective depths. Performance declines gradually as LP increases, followed by a sharp drop beyond a certain threshold. Specifically, this occurs after reducing effective depth by 9 layers for Qwen3 14B, by 7 layers for Llama2 7B and Qwen3 4B, and by 5 layers for Llama3.2 3B. These results indicate that larger models are more robust to the computational graph modifications from LP, suggesting that our approach is likely applicable to current commercial-scale LLMs used in major deployments.
It is worth noting that, unlike the other benchmarks, GSM-8K already drops severely in performance when applying low amounts of LP. Recent mechanistic interpretability research shows that LLMs have special circuitry for math operations, localized in a small set of parameters [46, 47]. [48] identify math specific parameters, and report a drop in accuracy of 17% when pruning them. We believe that the changes in the computational graph by the use of LP in some of the late layers of the LLM interfere with these fragile and sparse subnetworks, while leaving general language competence largely intact.
Impact on the inference speed. We run an ablation over several configurations and input sequence lengths on Figure 7 to test the speed on three different tasks: KV-Cache pre-filling, autoregressive generation up to the sequence length(with KV-Cache) and 1-token generation with a pre-filled KV-Cache of the corresponding sequence length. Our ablations show that the speed gain is directly proportional to the reduction of the effective depth of the model. For the effective depths of 25 ($\Delta=14$) in Llama 2 7B, we observe an average speed-up of 1.29x at the largest sequence length in the 1-token generation task. Likewise, for an effective depth of 23 ($\Delta=10$) in Llama 3.2 3B, we report a speed-up of 1.22x. For more aggressive parallelism, $\Delta=18$ and $\Delta=16$, we report a speed-up of 1.38x and 1.35x , at the expense of a large drop in ICL accuracy.
Fine-tuning for performance recovery. While LP provides speed improvements, associated architectural modifications may degrade model performance. To counteract this, we explored whether fine-tuning could effectively restore the original model’s capabilities. Specifically, we applied LP to layers 14–32 of Qwen3 4B, subsequently fine-tuning only these modified layers on randomly selected samples from the RedPajama training dataset [1]. Employing a batch size of 32, a linear learning rate schedule starting at $1e{-4}$, and the AdamW optimizer [49], we observed significant restoration of the model’s performance. Notably, GSM-8K accuracy improved from near-zero levels to more than half of the original model’s accuracy. It is possible that additional fine-tuning, or smarter tuning strategies, could yield further improvement, but resource constraints limited the scope of our experiments.
6 Limitations
The effectiveness of our approach exhibits notable variations across model scales. Smaller models show reduced benefits, likely due to their less sparse activation patterns and more tightly coupled layer dependencies. This degradation becomes more pronounced as the LP sequence increases, suggesting a practical upper limit to the number of layer pairs that can be effectively parallelized.
Finetuning. While some performance loss can be mitigated through fine-tuning, we were unable to fully recover the baseline model’s performance levels. This suggests fundamental trade-offs between computational efficiency and model capability that cannot be entirely eliminated through optimization.
Determining the ’true’ effective depth—the optimal configuration of parallel layer pairs—remains an open challenge as there is no theoretical framework for predicting the optimal grouping strategy.
These limitations highlight important directions for future research, particularly in developing more robust methods for determining optimal layer groupings and investigating the interplay between our approach and other efficiency-oriented techniques.
7 Conclusion
In this work, we presented Layer Parallelism, a novel approach that exploits independence patterns between transformer layers to optimize LLM inference. By restructuring the computational graph to enable parallel execution of consecutive layer pairs through tensor parallelism, we achieved substantial speed improvements without model retraining. We show that we can recover ICL accuracy by fine-tuning the parallelized models.
These results challenge the conventional view that transformer layers must process information strictly sequentially, suggesting instead that certain layers can operate independently without significant performance loss. From a practical standpoint, LP offers a straightforward approach to improve inference efficiency in production environments. Future work could focus on developing theoretical frameworks to predict optimal layer groupings, investigating interactions with other efficiency techniques such as quantization, and understanding the fundamental principles behind layer independence. Despite its limitations, LP represents a practical advancement in making LLM deployment more efficient and economically viable.
Impact Statement
This paper presents work whose goal is to improve inference serving of LLMs by trading accuracy. Well-intentioned users could find themselves receiving incorrect predictions that could lead to undesirable outcomes.
Acknowledgments
We thank Benjamin Rio for his insightful conversations and valuable input during the preparation of this work. His perspectives influenced several key aspects of the presentation.
Appendix A Ablation: Tokens per second
Figure 8: Tokens per second when completing the following inference tasks: KV Cache pre-filling for a given sequence length, autoregressive generation up to the indicated sequence length, and single token generation with a pre-filled KV Cache of the indicated sequence length. The baseline is the original model with all layers making use of Tensor Parallelism. The Parallel Layers number ($\Delta$) indicates how many layers have been merged using Layer Parallelism (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The number of tokens is computed as the sum of the input tokens and the output tokens for each forward pass.
Appendix B Generalization to multiple GPUs
[…]
Figure 9: Layer Parallelism in the case of parallelizing two layers over four accelerators. The stacked layers represent the tensor parallelism, and the colors indicate the processing of different previously contiguous layers. $Q,K,V\text{ and }att\in\mathbb{R}^{T\times\frac{2D}{g}}$, where $D$ is the feature dimension and $g$ is the total number of accelerators.
Layer Parallelism allows one to allocate $N\geq 1$ accelerators for each layer. The implementation remains the same, but now each layer is parallelized using tensor parallelism over its assigned accelerators. Note that both reduction operations (tensor parallel and layer parallel) are nicely executed with a single all-reduce call.
Appendix C Acceleration source
(a) Flame chart of running two standard Tensor-Parallel LLama decoder layers.
(b) Flame chart of running two Llama 3 decoder layers with our Layer Parallelism approach.
Figure 10: Comparison of Flame Graphs when running two consecutive Llama 3.2 3B decoder layers with vanilla tensor parallelism (Fig. 10(a)), and our Layer Parallelism approach (Fig. 10(b)). Note that the time axis scale is different between both graphs. These results were obtained on a workstation using x2 RTX 4090s.
Figure 10 illustrates flame graphs comparing two consecutive Llama 3.2 3B decoder layers using vanilla tensor parallelism and our LP approach. The profiling data summarized in Table 3 reveals that the primary source of acceleration in our LP method stems from reducing the total number of all reduce synchronization operations across GPUs. Specifically, the vanilla tensor parallel approach performs synchronization at every decoder layer, resulting in higher cumulative synchronization overhead. In contrast, our Layer Parallelism implementation runs pairs of layers simultaneously, effectively halving the number of synchronization points. This reduction in synchronization leads to a significant drop in synchronization time from 100.8ms to 50.7ms, directly contributing to the observed improvement in inference speed.
Additionally, Layer Parallelism enables fusion of certain computation kernels—particularly attention and MLP operations across parallelized layers—which further marginally reduces the computation time from 217ms to 208.7ms. Although these computational gains are modest compared to the savings achieved through fewer synchronization operations, kernel fusion further optimizes hardware utilization and enhances overall throughput.
Table 3: Profiling results comparing vanilla Tensor Parallel and Layer Parallel implementations on two consecutive Llama 3.2 3B decoder layers.
Approach Total Time (ms) Sync Time (ms) Computation Time (ms)
Tensor Parallel 317.8 100.8 217.0
Layer Parallel (Ours) 259.4 (x1.23) 50.7 (x1.99) 208.7 (x1.04)